OpenStudy (anonymous):

use the information on the state park sign shown on the right. suppose you bicycle from the park sign to the picnic parking, then bicycle back to lake, and finally bicycle to the beach parking. how far will you have traveled? write an exact answer in simplest form. Lake:1/2 mi cafe:1 3/4mi picnic parking:3 3/4mi beach parking:3 7/8mi

OpenStudy (ash2326):

\[Lake\ = \frac{1}{2} mi\] \[Cafe\ = 1\frac{3}{4} mi\] \[Picnic Parking\ = 3\frac{3}{4} mi\] \[Beach Parking\ = 3\frac{7}{8} mi\]

OpenStudy (ash2326):

We bicycle in the following way \[Park\ Sign\ \to\ Picnic\ Parking\ \to\ Park\ Sign\ \to\ lake\ \to\] \[Park\ Sign\ \to\ Beach\ Parking\ \]

OpenStudy (ash2326):

\[3\frac{3}{4}+3\frac{3}{4}+\frac{1}{2}+\frac{1}{2}+3\frac{7}{8}\] \[3+ \frac{3}{4}+3+ \frac{3}{4}+\frac{1}{2}+3 +\frac{7}{8}\] In the denominators we have 2, 4 and 8. The LCM of these three is 8 Can you solve now @Jyako90?

OpenStudy (anonymous):

6/8 6/8 4/8

OpenStudy (ash2326):

@Jyako90 If you need any help my friend @rebeccaskell94 will help you. I gotta run Sorry:(

OpenStudy (anonymous):

aww ;/

OpenStudy (ash2326):

You'll get \[11 \frac{7}{8}\]

OpenStudy (ash2326):

Try to figure how you get this. Bye:)

OpenStudy (anonymous):

thats the answer

OpenStudy (ash2326):

But do figure how you get this. I'll check it when I come back

OpenStudy (anonymous):

okay let me see

OpenStudy (anonymous):

okay so like basically its kinda like later on in algebra where what you do to one side you have to do to another. in this case if you multiply the denominator by say 4 you have oto muliply the numerator by 4 also. sooo let me do a draw and Ill post that

OpenStudy (anonymous):

ok so the LDC is 8

OpenStudy (anonymous):

yes ma'am all of the denominators can be multiplied to 8

OpenStudy (anonymous):

oh wait...we need to make the mixed numbers improper fractions first. So now we have: 15/4 + 15/4 + 1/2 + 1/2 + 31/8 made into LCD we have 30/8 + 30/8+ 4/8 + 4/8 + 31/8 :)

OpenStudy (anonymous):

which is 99/8 or 12 3/8

OpenStudy (anonymous):

ok 95/8= 11 7/8

OpenStudy (anonymous):

she said the answer is 11 7/8?

OpenStudy (anonymous):

I thnk he most of added wrong... (aSh is most def a guy) lol

OpenStudy (anonymous):

I'll have him double check it though :)

OpenStudy (anonymous):

haha ok so

OpenStudy (anonymous):

so is that it 12 3/8

OpenStudy (anonymous):

I'm pretty sure it is :)

OpenStudy (anonymous):

okay thanks :)